# Robust Barycenters of Persistence Diagrams

### v1

@@ IV-C Computation time comparison
TABLE I: Running times (in seconds) for computing a $W_{2}$ barycenter.
Method $m=4$ $m=6$
Arithmetic Mean 1.9 24.9
$\mathfrak{b}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2}}$ (Pytorch) 6.3 29.0
In this section, we compare the time needed to compute a barycenter, with regard to the $W_{2}$ metric, using the arithmetic mean between points of $\mathbb{R}^{2}$, and the time when computing $\mathfrak{b}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2}}$. Our experiment consists in computing a barycenter of persistence diagrams for the Isabel ensemble (Sec. II-B), where each diagram is threshold by persistence to feature only $\sim 100$ pairs and where $T$ is set to $5$. We report the resulting running times in Tab. I. When considering $m=4$ input diagrams, the computation based on the arithmetic mean is 3 times faster than the one based on $\mathfrak{b}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2}}$. However, for a larger ensemble ($m=6$), the difference is reduced to $4$ seconds, yielding compatible run times between the two approaches.
@@ V Conclusion
In this paper, we showcased the utility of a method for computing a robust Wasserstein barycenters of persistence diagrams. Specifically, we adapted a recent fixed-point method algorithm [74] to the case of persistence diagrams. We first gave a reminder on this fixed point method framework to compute this robust barycenter. We also gave a formal proof of the necessary hypothesis for the convergence of this method in the Appendix. Then we presented two applications of this robust barycenter to clustering and dictionary encoding of persistence diagrams in the presence of outliers. We believe that our work on the robustness of Wasserstein barycenters is a useful step toward improving their applicability in the analysis of real-life ensembles of persistence diagrams.
@@ Acknowledgments
Appendix A Sufficient Conditions for Assumption 1

### v2

@@ IV-C Computation time comparison
TABLE I: Running times (in seconds) for computing a $W_{q}$ barycenter.
Method $m=4$ $m=6$ $m=8$ $m=10$ $m=12$
Arithmetic Mean 0.1 0.3 4.5 36.7 100.1
$\mathfrak{b}_{{2}}$ (Pytorch) 5.4 5.0 9.7 47.3 147.7
$\mathfrak{b}_{{1.8}}$ (Pytorch) 5.1 5.0 9.9 50.1 135.5
$\mathfrak{b}_{{1.6}}$ (Pytorch) 5.0 5.4 9.4 54.1 149.2
$\mathfrak{b}_{{1.4}}$ (Pytorch) 4.4 6.5 10.6 56.0 147.6
$\mathfrak{b}_{{1.2}}$ (Pytorch) 5.2 7.6 11.6 57.6 152.7
In this section, we compare the time needed to compute a barycenter, with regard to the $W_{2}$ metric, using the arithmetic mean between points of $\mathbb{R}^{2}$, and the time when computing $\mathfrak{b}_{2}$. Our experiment consists in computing a barycenter of persistence diagrams for the Isabel ensemble (Sec. II-B), where each diagram is thresholded by persistence to feature only $\sim 20$ pairs and where $T$ is set to $5$ for this experiment. We report the resulting running times in Tab. I. When considering $m=4,6$ input diagrams, the computation based on the arithmetic mean is 50 times faster than the one based on $\mathfrak{b}_{{2}}$. However, for a larger ensemble ($m=8,10,12$), the difference is reduced. This is because diagrams need to be augmented prior to computing matchings between them, and that, consequently, the size of the barycenter increases significantly when $m$ increases, yielding an assignment time (cubic complexity) being prevalent over the update time (see Alg. 1 and Sec. III-A).
@@ V Conclusion
In this paper, we showcased the utility of a method for computing a robust Wasserstein barycenters of persistence diagrams. Specifically, we adapted a recent fixed-point method algorithm [76] to the case of persistence diagrams. We first gave a reminder on this fixed point method framework to compute this robust barycenter, and then gave insights on the choice of $q$ depending on the characteristic one wants from the barycenter. We also gave a formal proof of the necessary hypothesis for the convergence of this method in the Appendix. Then we presented two applications of this robust barycenter to clustering and dictionary encoding of persistence diagrams in the presence of outliers. We also noticed that, across all our experiments, taking $q\in[1.2,1.4]$ yields the best results. We believe that our work on the robustness of Wasserstein barycenters is a useful step toward improving their applicability in the analysis of real-life ensembles of persistence diagrams.
@@ Acknowledgments
Appendix A Parameters for Adam Optimizer
Regarding the optimizer used to update the barycenter in the algorithm, we used the following parameters: a learning rate $\ell r=1$, and the optimization stops if either the maximum number of iterations $T^{\prime}=250$ is reached or two iterates of the energy are close (under a threshold of $10^{-15}$).
Appendix B Sufficient Conditions for Assumption 1